Profit Margin, Investment Turnover, and return on investment The condensed income statement for the Consumer Products...

  1. Profit Margin, Investment Turnover, and return on investment

    The condensed income statement for the Consumer Products Division of Fargo Industries Inc. is as follows (assuming no service department charges):

    Sales $1,146,000
    Cost of goods sold 515,700
    Gross profit $630,300
    Administrative expenses 343,800
    Income from operations $286,500

    The manager of the Consumer Products Division is considering ways to increase the return on investment.

    a. Using the DuPont formula for return on investment, determine the profit margin, investment turnover, and return on investment of the Consumer Products Division, assuming that $1,910,000 of assets have been invested in the Consumer Products Division. Round the investment turnover to one decimal place.

    Profit margin %
    Investment turnover
    Rate of return on investment %

    b. If expenses could be reduced by $57,300 without decreasing sales, what would be the impact on the profit margin, investment turnover, and return on investment for the Consumer Products Division? Round the investment turnover to one decimal place.

    Profit margin %
    Investment turnover
    Rate of return on investment %

Check My Work

  • Previous
  • Next
Answer #1

a) Profit margin = Income from operation/Sales = 286500/1146000 = 25%

Investment turnover = Sales/Average operating assets = 1146000/1910000 = 0.6

Rate of return on investment = Profit margin*Investment turnover = 25*.6 = 15%

b) Profit margin = Income from operation/Sales = 343800/1146000 = 30%

Investment turnover = Sales/Average operating assets = 1146000/1910000 = 0.6

Rate of return on investment = Profit margin*Investment turnover = 30*.6 = 18%